# Phantom Menace: Exploring and Enhancing the Robustness of VLA Models against Physical Sensor Attacks

### v1

@@ Abstract
Vision-Language-Action (VLA) models revolutionize robotic systems by enabling end-to-end perception-to-action pipelines that integrate multiple sensory modalities, such as visual signals processed by cameras and auditory signals captured by microphones. This multi-modality integration allows VLA models to interpret complex, real-world environments using diverse sensor data streams. Given the fact that VLA-based systems heavily rely on the sensory input, the security of VLA models against physical-world sensor attacks remains critically underexplored.
To address this gap, we present the first systematic study of physical sensor attacks against VLAs, quantifying the influence of sensor attacks and investigating the defenses for VLA models. We introduce a novel “Real-Sim-Real” framework that automatically simulates physics-based sensor attack vectors, including six attacks targeting cameras and two targeting microphones, and validates them on real robotic systems. Through large-scale evaluations across various VLA architectures and tasks under varying attack parameters, we demonstrate significant vulnerabilities, with susceptibility patterns that reveal critical dependencies on task types and model designs. We further develop an adversarial-training-based defense that enhances VLA robustness against out-of-distribution physical perturbations caused by sensor attacks while preserving model performance. Our findings expose an urgent need for standardized robustness benchmarks and mitigation strategies to secure VLA deployments in safety-critical environments.
@@ Introduction
Figure 1: Exploring and enhancing the robustness of VLA models against physical sensor attacks. We validate that VLA models are vulnerable to physical sensor attacks during their interaction with visual, audio signals and could misbehave in real-world applications. Attackers can induce four serious consequence by injecting laser, electromagnetic, and ultrasound signals into cameras and microphones. We present an automatic “Real-Sim-Real” framework to quantify the influence and propose the defense for VLA models against such physical sensor attacks.
@@
VLA-based systems heavily rely on sensory input, making their robustness and security in physical interactions increasingly important and timely concerns. However, in the current research landscape of VLA security, prior work (krzysztof2025adversarial; wang2024exploring; zhou2025badvla; cheng2024manipulation) has primarily focused on digitally manipulating the inputs of VLA models, rather than using physical signals. Typical attacks involve directly modifying images or text to generate adversarial input. While these attacks can be efficient, they may not fully reflect the unique characteristics inherent in physical-world interactions. Consequently, such attacks may fail to comprehensively capture the vulnerabilities that arise during real-world deployment scenarios.
@@
  • How to defend such physical sensor attacks and enhance the robustness of VLA modes?
To answer the above questions, we present an automatic “Real-Sim-Real” framework (as shown in Figure 1) to validate VLA model robustness against physical sensor attacks effectively and realistically both in the simulator and real world. First, we systematically review existing physical sensor attack techniques and select eight representative examples from top-tier security conferences: six targeting cameras and two targeting microphones. These attacks interfere with VLA models by injecting physical signals—such as ultrasound, laser, or electromagnetic waves—into sensors. Unlike passive attacks that rely on physical adversarial examples (brown2017adversarial), attackers can actively control the initiation and termination of these sensor attacks, thereby achieving stealthiness in real-world scenarios. Next, we develop high-fidelity digital simulations of these physical sensor attacks based on their underlying physical principles and patterns observed in actual attacks. We define three levels of attack intensity, namely, strong, medium, and weak to achieve different attack consequences. Finally, we conduct large-scale robustness evaluations of four VLA models across four datasets within this simulation environment.
@@
  • We propose a “Real-Sim-Real” framework to validate VLA model robustness against physical sensor attacks in a systematical and realistic way. This framework effectively bridges the gap between purely digital attack simulations and resource-intensive physical experiments.
@@
With the rapid advancement of VLA models, their security has garnered significant attention. Yet, existing studies have primarily centered on vulnerability exploration in digital domain. Specifically, the RoboticAttack (wang2024exploring) framework introduces adversarial patch attacks targeting image inputs. Robotgcg (krzysztof2025adversarial) applies LLM jailbreak attacks in the text modality. BadVLA (zhou2025badvla) framework represents the VLA backdoor attacks. Different from this work, we are the first to explore physical sensor attacks on VLA models.
@@
Current VLA robustness evaluation studies focus on scenario generalization and are primarily conducted in simulation environments. For instance, PVEP (cheng2024manipulation) evaluates robustness under blurring, Gaussian noise, brighter, and darker conditions. VLATest (wang2024towards) examines the effects of obstacles, lighting conditions, camera poses, and unseen objects. Unlike these studies that explore in-distribution robustness, we investigate the out-of-distribution robustness of VLA models from the sensor attack perspective, reflecting practical risks of VLA models in real-world deployments.
@@
A VLA model receives visual images and audio instructions through cameras and microphones to generate robot actions end-to-end. As illustrated in Figure 2, visual inputs typically consist of RGB images captured by multiple cameras (e.g., full camera, wrist camera, etc.), whereas audio instructions are collected via microphones and subsequently converted to text using Automatic Speech Recognition (ASR). Generally, a VLA comprises two primary components: a Vision-Language Model (VLM) and an action decoder. The VLM (driess2023palm; bai2025qwen2; beyer2024paligemma; alayrac2022flamingo; liu2023visual; peng2023kosmos) pre-trained on large-scale multimodal internet datasets extracts action tokens from the sensor perception data. Based on these action tokens, the action decoder generates corresponding robot actions to interact with the environment through physical systems, such as robotic arms (zhong2025survey).
Vision-Language Model. The VLM typically consists of visual encoders and a Large Language Model (LLM) backbone. To effectively extract high-level visual features, pretrained visual encoders, such as CNN (he2016deep), ViT (dosovitskiy2020image), SigLIP (zhai2023sigmoid), and DinoV2 (oquab2023dinov2), are commonly utilized. The LLM backbone is generally a small parameter (always under 7B), such as Llama2 (touvron2023llama), Palm (chowdhery2023palm), T5 (raffel2020exploring), or Gemma (team2024gemma), which convert visual embeddings and textual embeddings to action tokens.
Action Decoder. Action decoders can be categorized into multiple types based on how they convert action tokens to physical robot actions. Direct decoding treats LLM output textual tokens as action tokens, generating discrete robot actions token-by-token autoregressively (brohan2023rt; kim2024openvla). MLP mapping converts LLM hidden states into robot actions using a simple MLP (kim2025fine; zheng2025universal). Diffusion decoding gradually denoises random noisy action sequences to reconstruct robot actions under the condition of LLM feature vectors (wen2025dexvla; chi2023diffusion). Flow matching decoding directly matches the velocity field from the current distribution to the target action distribution by learning mappings from control signals to actions (black2024pi_0; shukor2025smolvla).
@@
Camera and attacks. A camera primarily consists of a light-sensitive transducer, signal processing circuitry, and an image signal processor (ISP). The transducer converts optical signals into electrical signals, which are then denoised, amplified, and digitized. The ISP further enhances the digital image through various compensation and correction processes. Some cameras also integrate an Inertial Measurement Unit (IMU) for image stabilization. The above process can be disrupted by lasers (yan2022rolling) and acoustic signals (ji2021poltergeist), compromising measurement accuracy.
Microphone and attacks. A microphone mainly consists of an acoustic transducer and a signal processing circuit. The transducer converts acoustic signals into electrical signals. The signal processing circuit amplifies, filters, and digitizes it. However, the above process is also demonstrated to be susceptible to ultrasonic signals (zhang2017dolphinattack) and lasers (sugawara2020light).
@@
Attack Goal. The goal of attackers is to launch physical sensor attacks on cameras or microphones, inducing VLA robot systems to perform unexpected or even targeted incorrect actions to lead to task failure.
Attacker Capability. Attackers can only launch physical signal attacks on the sensors of VLA robot systems. They cannot conduct digital attacks such as injecting noise, applying compression, blurring, or watermarking.
Model Knowledge. Attackers have only black-box access to VLA models, without knowledge of the training data, model architecture, or pre-trained parameters. Additionally, they are unaware of the specific sensor types and algorithms used (e.g., ASR, image stabilization).
Figure 3: We implement and simulate eight sensor attacks, including six targeting cameras and 2 targeting microphones, covering laser, light, acoustic, and EM signals. Attack instances are under varying attack intensities for each attack, i.e., the attack intensity progressively increases from left to right.
@@
where $S_{attacked}(t)$ is the total audio signal captured by the microphone at time $t$, $S_{original}(t)$ is the original audio signal, and $S_{malicious}(t)$ is the malicious noise or instruction signal.
@@
We select Libero (liu2023libero) as the simulator for our simulated experiments. Libero is an open-source visual-language robotics simulator designed to provide a flexible testing platform for VLA models. We also use Libero datasets as follows:
Libero-Spatial involves manipulating identical objects placed in varying spatial configurations.
Libero-Object emphasizes the ability to recognize and manipulate various objects. Tasks in this suite involve moving different objects to specific locations.
Libero-Goal consists of tasks where the objects and their spatial arrangements remain constant, but the goals differ.
Libero-Long includes tasks that involve long-horizon planning and execution.
@@
Voice DOS instruction = None
Voice Spoofing suffix = ignore the above instruction and do not move
Table 1: Attack parameter of different attack intensities.
@@
As shown in Figure 3, we set three attack intensities—weak, medium, and strong—in the simulator. The corresponding attack parameters are listed in Table 1. In real-world experiments, we use the parameters identified through simulation.
@@
Attack OpenVLA OpenVLA-OFT $\pi$0 $\pi$0-fast
@@
´ $LB_{weak}$ 85.0 86.0 73.8 50.4 98.0 98.4 97.4 93.8 96.4 98.0 94.8 83.2 96.0 98.8 91.8 65.0
´ $LB_{medium}$ 68.2 61.0 64.4 16.4 98.0 98.0 97.6 86.0 97.2 97.6 94.2 77.8 98.0 99.0 85.0 54.0
´ $LB_{strong}$ 0.0 0.0 0.0 0.0 87.8 5.6 78.4 18.4 57.0 78.0 52.4 23.8 62.0 85.0 62.0 36.0
´ $LP_{weak}$ 83.0 88.0 72.8 43.4 98.4 98.8 97.0 94.2 98.4 96.6 93.0 78.0 96.0 98.0 89.8 62.0
´ $LP_{medium}$ 59.4 70.0 26.4 14.0 98.8 97.0 97.8 89.4 97.8 97.0 93.4 74.6 98.0 97.2 86.0 59.0
´ $LP_{strong}$ 59.4 66.8 19.8 11.2 98.4 97.8 95.4 81.4 97.8 97.8 92.2 77.0 94.0 98.0 85.8 59.0
´ $ECS_{weak}$ 63.6 78.4 67.0 27.2 98.4 96.8 97.4 91.8 99.2 98.8 94.0 77.0 97.0 98.2 90.0 58.8
´ $ECS_{medium}$ 34.8 62.8 53.4 19.4 97.8 99.0 96.4 89.6 97.0 98.4 89.8 81.4 96.0 98.4 85.0 55.0
´ $ECS_{strong}$ 45.0 59.0 65.2 16.0 98.2 98.8 97.0 90.2 97.8 99.0 94.6 81.4 96.0 98.6 88.0 53.6
´ $ET_{weak}$ 24.0 20.4 25.2 2.0 92.2 92.4 95.2 72.4 96.2 95.6 90.0 76.2 95.0 96.4 88.0 58.0
´ $ET_{medium}$ 4.6 0.6 11.6 0.0 89.8 74.0 89.6 45.0 96.0 96.0 82.2 60.8 99.0 96.0 78.0 50.2
´ $ET_{strong}$ 0.4 0.0 8.4 0.0 96.4 54.8 87.4 26.8 95.2 94.8 70.0 44.8 95.0 93.0 69.0 48.0
´ $LCS_{weak}$ 72.2 65.2 57.0 12.2 97.2 98.4 97.6 87.2 97.4 98.6 93.6 81.4 96.0 98.2 85.0 55.0
´ $LCS_{medium}$ 44.6 34.0 17.4 1.6 97.4 97.8 97.4 68.4 97.4 99.4 88.2 75.6 97.0 94.0 75.8 51.0
´ $LCS_{strong}$ 11.8 2.0 9.8 0.0 95.4 94.2 90.4 51.8 94.0 96.0 74.0 40.4 94.0 93.0 72.0 51.0
´ $UB_{weak}$ 3.4 1.8 10.6 3.8 98.2 96.4 98.4 89.6 97.8 97.4 93.2 82.8 94.0 96.0 79.0 53.0
´ $UB_{medium}$ 0.2 0.0 0.0 0.0 96.8 51.0 89.4 36.2 96.6 96.8 88.4 73.4 93.0 93.2 74.0 55.0
´ $UB_{strong}$ 0.0 0.0 0.0 0.0 90.8 9.4 68.0 10.2 82.0 83.4 49.4 30.8 82.0 96.0 52.0 41.4
@@
Table 2: Robustness of VLA models in the simulator. ´: Camera; : Microphone; LB: Laser Blinding; LP: Light Projection; ECS: EM Color Strip; ET: EM Truncation; LCS: Laser Color Strip; UB: Ultrasound Blur; VD: Voice DoS; VS: Voice Spoofing.
@@
Attack OpenVLA OpenVLA-OFT $\pi$0 $\pi$0-fast
@@
´ AT $LB_{medium}$ 78.8 84.0 76.6 42.2 98.6 99.0 97.0 93.2 98.4 97.6 93.8 77.2 97.2 97.8 86.2 57.2
´ AT $LP_{medium}$ 66.2 74.4 28.4 11.8 98.6 97.6 96.8 91.6 97.8 96.2 94.4 78.6 97.4 97.8 86.0 58.2
´ AT $ECS_{medium}$ 72.0 76.0 63.6 34.8 98.2 98.0 96.8 92.4 98.6 96.4 94.4 79.6 95.6 98.6 84.2 58.0
´ AT $ET_{medium}$ 35.6 3.4 23.4 0.0 99.4 96.6 97.0 79.8 98.4 97.6 92.4 76.4 96.4 97.6 82.6 52.4
´ AT $LCS_{medium}$ 82.0 85.2 69.2 36.2 98.0 98.4 96.8 90.0 98.4 98.4 91.6 78.0 97.0 98.0 84.6 59.4
´ AT $UB_{medium}$ 56.8 66.6 46.2 8.0 97.6 99.0 97.2 92.2 97.6 97.4 94.4 82.6 97.8 98.6 83.2 57.8
@@
Answer 1 Physical sensor attacks succeed on VLA models.
Performance without sensor attacks. We first evaluate the performance of four VLA models after fine-tuning on Libero datasets. As shown in Table 2, these models demonstrate strong performance, achieving up to 90% TSR on simple tasks such as Libero-Spatial and Libero-Object. For long-horizon tasks, OpenVLA-OFT maintains good performance. These results indicate that VLA models possess robust task execution capabilities across diverse manipulation scenarios.
Performance against simulated sensor attacks. Then, we evaluate the performance of four VLA models against simulated eight sensor attacks. As shown in Table 2, all VLA models clearly exhibit vulnerability to sensor attacks. The degree of performance degradation differs significantly depending on the specific VLA architecture, attack type, and attack intensity. In many scenarios, particularly under strong attacks or during long-horizon tasks, model performance collapses catastrophically.
@@
Answer 2.1 The impacts of sensor attacks on VLA models vary.
@@
Camera attacks, including Laser Blinding (LB), EM Truncation (ET), and Ultrasound Blur (UB), highly disrupt the visual information of VLA models, especially under medium to strong intensity settings. The absence of object locations and class identities leads to severe task failures, resulting in unexpected actions or potentially harmful operations. In contrast, other camera attacks, such as Light Projection (LP), Laser Color Strip (LCS), and EM Color Strip (ECS), are comparatively less severe. These attacks interfere with the models’ attention by injecting perturbation rather than destroying critical visual features. Consequently, the primary objects and intended goals typically remain discernible, resulting in less degradation in TSR. Voice DoS attacks adversely affect VLA model performance, with the severity primarily determined by dataset characteristics. In the Libero-Goal dataset, where robots must execute varying instructions within fixed scenes, the absence of voice instructions leaves VLA models unable to determine appropriate actions. Conversely, datasets featuring unique scene-instruction correspondences (e.g., Libero-Spatial, Libero-Object, and Libero-Long) enable VLA models to infer correct instructions from visual context alone, thereby reducing the attack’s effectiveness. The effectiveness of the voice spoofing attack is strongly correlated with the target VLA model’s semantic understanding and instruction-following capabilities. Compared to pi0 and pi0-fast, OpenVLA and OpenVLA-OFT employ LLM backbones, making them more vulnerable to malicious instruction injection. OpenVLA-OFT incorporates a FiLM module that leverages task-specific language embeddings to modulate visual features, thereby enhancing instruction-following capabilities beyond OpenVLA. Consequently, OpenVLA-OFT exhibits the most substantial performance degradation under this attack.
Answer 2.2 VLA models exhibit different robustness against sensor attacks.
@@
Attack OpenVLA OpenVLA-OFT pi0 pi0-fast
None Attack 5/10 8/10 10/10 10/10
´ Laser Blinding 0/10 0/10 0/10 0/10
´ Light Projection 0/10 1/10 1/10 0/10
´ EM Color Strip 0/10 0/10 6/10 0/10
´ EM Truncation 0/10 0/10 0/10 0/10
´ Laser Color Strip 0/10 0/10 3/10 0/10
´ Ultrasound Blur 0/10 0/10 1/10 0/10
@@
We first evaluate the performance of VLA models without sensor attacks to establish baseline model availability. Subsequently, we inject malicious signals into sensors in real-world scenarios using attack parameters identified through simulation. Table 4 presents the results, demonstrating strong alignment between physical and simulation conclusions, thereby validating the effectiveness of the attack parameters searched in the simulation. As shown in Figure 5, the attacks could induce four distinct consequences: 1) previously closed grippers unexpectedly release, causing objects to fall and sustain damage; 2) robotic arms or grippers collide with objects or environmental structures, damaging either the objects or the grippers themselves; 3) robotic arms grasp incorrect objects, preventing successful task completion; and 4) robotic arms exhibit erratic movement patterns, resulting in chaos and energy waste. Demo videos could be found in the supplementary material.

### v2

@@ Abstract
Vision-Language-Action (VLA) models revolutionize robotic systems by enabling end-to-end perception-to-action pipelines that integrate multiple sensory modalities, such as visual signals processed by cameras and auditory signals captured by microphones. This multi-modality integration allows VLA models to interpret complex, real-world environments using diverse sensor data streams. Given the fact that VLA-based systems heavily rely on the sensory input, the security of VLA models against physical-world sensor attacks remains critically underexplored. To address this gap, we present the first systematic study of physical sensor attacks against VLAs, quantifying the influence of sensor attacks and investigating the defenses for VLA models. We introduce a novel “Real-Sim-Real” framework that automatically simulates physics-based sensor attack vectors, including six attacks targeting cameras and two targeting microphones, and validates them on real robotic systems. Through large-scale evaluations across various VLA architectures and tasks under varying attack parameters, we demonstrate significant vulnerabilities, with susceptibility patterns that reveal critical dependencies on task types and model designs. We further develop an adversarial-training-based defense that enhances VLA robustness against out-of-distribution physical perturbations caused by sensor attacks while preserving model performance. Our findings expose an urgent need for standardized robustness benchmarks and mitigation strategies to secure VLA deployments.
@@ Introduction
Figure 1: Overview of the “Real-Sim-Real” framework. We demonstrate that VLA models are vulnerable to physical sensor attacks, where attackers inject malicious signals (e.g., laser, electromagnetic interference, ultrasound) into cameras and microphones, leading to severe consequences in real-world deployments. Our framework automatically evaluates these physical attack vectors to quantify their impact and we propose defenses for enhancing VLA robustness.
@@
VLA-based systems heavily rely on sensory input, making their robustness and security in physical interactions increasingly important and timely concerns. However, in the current research landscape of VLA security, prior work (krzysztof2025adversarial; wang2024exploring; zhou2025badvla; cheng2024manipulation) has primarily focused on digitally manipulating the inputs of VLA models, rather than using physical signals. Typical attacks involve directly modifying images or text to generate adversarial input. While these attacks can be efficient, they may not fully reflect the unique characteristics inherent in physical-world interactions. Consequently, such attacks may fail to comprehensively capture the vulnerabilities that arise during real-world deployments.
@@
  • How to defend against such physical sensor attacks and enhance the robustness of VLA modes?
To answer the above questions, we present an automatic “Real-Sim-Real” framework (as shown in Figure 1) to validate VLA model robustness against physical sensor attacks effectively and realistically both in the simulator and real world. First, we systematically review existing physical sensor attack techniques and select eight representative examples from top-tier security conferences: six targeting cameras and two targeting microphones. These attacks interfere with VLA models by injecting physical signals—such as ultrasound, laser, or electromagnetic waves—into sensors. Unlike passive attacks that rely on physical adversarial examples (brown2017adversarial), attackers can actively control the initiation and termination of these sensor attacks, thereby achieving stealthiness in real-world scenarios. Next, we develop high-fidelity digital simulations of these physical sensor attacks based on their underlying physical principles and patterns observed in actual attacks. We define three levels of attack intensity, namely, strong, medium, and weak, to achieve different attack consequences. Finally, we conduct large-scale robustness evaluations of four VLA models across four datasets within this simulation environment.
@@
  • We propose a “Real-Sim-Real” framework to validate VLA model robustness against physical sensor attacks in a systematic and realistic way. This framework effectively bridges the gap between purely digital attack simulations and resource-intensive physical experiments.
@@
With the rapid advancement of VLA models, their security has garnered significant attention. Yet, existing studies have primarily centered on vulnerability exploration in the digital domain. Specifically, the RoboticAttack (wang2024exploring) framework introduces adversarial patch attacks targeting image inputs. Robotgcg (krzysztof2025adversarial) applies LLM jailbreak attacks in the text modality. BadVLA (zhou2025badvla) framework represents the VLA backdoor attacks. Different from this work, we are the first to explore physical sensor attacks on VLA models.
@@
Current VLA robustness evaluation studies focus on scenario generalization and are primarily conducted in simulation environments. For instance, PVEP (cheng2024manipulation) evaluates robustness under blurring, Gaussian noise, and different lighting conditions. VLATest (wang2024towards) examines the effects of obstacles, lighting conditions, camera poses, and unseen objects. Unlike these studies that explore in-distribution robustness, we investigate the out-of-distribution robustness of VLA models from the sensor attack perspective, reflecting practical risks of VLA models in real-world deployments.
@@
A VLA model receives visual images and audio instructions through cameras and microphones to generate robot actions end-to-end. As illustrated in Figure 2, visual inputs typically consist of RGB images captured by multiple cameras (e.g., full camera, wrist camera, etc.), whereas audio instructions are collected via microphones and subsequently converted to text using Automatic Speech Recognition (ASR). Generally, a VLA comprises two primary components: a Vision-Language-Model (VLM) and an action decoder. The VLM (driess2023palm; bai2025qwen2; beyer2024paligemma; alayrac2022flamingo; liu2023visual; peng2023kosmos) pre-trained on large-scale multimodal internet datasets, extracts action tokens from the sensor perception data. Based on these action tokens, the action decoder generates corresponding robot actions to interact with the environment through physical systems (zhong2025survey).
Vision-Language Model.
The VLM typically consists of visual encoders and a Large Language Model (LLM) backbone. To effectively extract high-level visual features, pretrained visual encoders, such as CNN (he2016deep), ViT (dosovitskiy2020image), SigLIP (zhai2023sigmoid), and DinoV2 (oquab2023dinov2), are commonly utilized. The LLM backbone is generally a small parameter (always under 7B), such as Llama2 (touvron2023llama), Palm (chowdhery2023palm), T5 (raffel2020exploring), or Gemma (team2024gemma), which convert visual embeddings and textual embeddings to action tokens.
Action Decoder.
Action decoders can be categorized into multiple types based on how they convert action tokens to physical robot actions. Direct decoding treats LLM output textual tokens as action tokens, generating discrete robot actions token-by-token autoregressively (brohan2023rt; kim2024openvla). MLP mapping converts LLM hidden states into robot actions using a simple MLP (kim2025fine; zheng2025universal). Diffusion decoding gradually denoises random noisy action sequences to reconstruct robot actions under the condition of LLM feature vectors (wen2025dexvla; chi2023diffusion). Flow matching decoding matches the velocity field from the current distribution to the target distribution by learning mappings from control signals to actions (black2024pi_0; shukor2025smolvla).
@@
Camera and attacks.
A camera primarily consists of a light-sensitive transducer, signal processing circuitry, and an image signal processor (ISP). The transducer converts optical signals into electrical signals, which are then denoised, amplified, and digitized. The ISP further enhances the digital image through various compensation and correction processes. Some cameras also integrate an Inertial Measurement Unit (IMU) for image stabilization. The above process can be disrupted by lasers (yan2022rolling) and acoustic signals (ji2021poltergeist), compromising measurement accuracy.
Microphone and attacks.
A microphone mainly consists of an acoustic transducer and a signal processing circuit. The transducer converts acoustic signals into electrical signals. The signal processing circuit amplifies, filters, and digitizes it. However, the above process is also demonstrated to be susceptible to ultrasonic signals (zhang2017dolphinattack) and lasers (sugawara2020light).
Methodology
@@
Attack Goal.
The goal of attackers is to launch physical sensor attacks on cameras or microphones, inducing VLA robot systems to perform unexpected or even targeted incorrect actions to lead to task failure.
Attacker Capability.
Attackers can only launch physical signal attacks on the sensors of VLA robot systems. They cannot conduct digital attacks such as injecting noise, applying compression, blurring, or watermarking.
Model Knowledge.
Attackers have only black-box access to VLA models, without knowledge of the training data, model architecture, or pre-trained parameters. Additionally, they are unaware of the specific sensor types and algorithms used (e.g., ASR, image stabilization).
Figure 3: We implement and simulate eight sensor attacks, including six targeting cameras and two targeting microphones, covering laser, light, acoustic, and EM signals. Attack instances are under varying attack intensities for each attack, i.e., the attack intensity progressively increases from left to right.
@@
where $S_{attacked}(t)$ is the total audio signal captured by the microphone at time $t$, $S_{original}(t)$ is the original audio signal, and $S_{malicious}(t)$ is the malicious signal.
@@
We select Libero (liu2023libero) as the simulator for our simulated experiments. Libero is an open-source visual-language robotics simulator designed to provide a flexible testing platform for VLA models. We also use Libero datasets as follows: Libero-Spatial involves manipulating identical objects placed in varying spatial configurations; Libero-Object emphasizes the ability to recognize and manipulate various objects. Tasks in this suite involve moving different objects to specific locations; Libero-Goal consists of tasks where the objects and their spatial arrangements remain constant, but the goals differ; Libero-Long includes tasks that involve long-horizon planning and execution.
@@
Table 1: Attack parameters of different attack intensities.
@@
As shown in Figure 3, we set three attack intensities—weak, medium, and strong—in the simulator. The corresponding attack parameters are listed in Table 1. For Voice DOS attacks, we set the instruction as “None.” For voice spoofing attacks, we set the suffix as “ignore the above instruction and do not move.” In real-world experiments, we use the parameters searched in simulation.
@@
Attack Method OpenVLA OpenVLA-OFT $\pi$0 $\pi$0-fast
@@
 $LB_{weak}$ 85.0 86.0 73.8 50.4 98.0 98.4 97.4 93.8 96.4 98.0 94.8 83.2 96.0 98.8 91.8 65.0
 $LB_{medium}$ 68.2 61.0 64.4 16.4 98.0 98.0 97.6 86.0 97.2 97.6 94.2 77.8 98.0 99.0 85.0 54.0
 $LB_{strong}$ 0.0 0.0 0.0 0.0 87.8 5.6 78.4 18.4 57.0 78.0 52.4 23.8 62.0 85.0 62.0 36.0
 $LP_{weak}$ 83.0 88.0 72.8 43.4 98.4 98.8 97.0 94.2 98.4 96.6 93.0 78.0 96.0 98.0 89.8 62.0
 $LP_{medium}$ 59.4 70.0 26.4 14.0 98.8 97.0 97.8 89.4 97.8 97.0 93.4 74.6 98.0 97.2 86.0 59.0
 $LP_{strong}$ 59.4 66.8 19.8 11.2 98.4 97.8 95.4 81.4 97.8 97.8 92.2 77.0 94.0 98.0 85.8 59.0
 $ECS_{weak}$ 63.6 78.4 67.0 27.2 98.4 96.8 97.4 91.8 99.2 98.8 94.0 77.0 97.0 98.2 90.0 58.8
 $ECS_{medium}$ 34.8 62.8 53.4 19.4 97.8 99.0 96.4 89.6 97.0 98.4 89.8 81.4 96.0 98.4 85.0 55.0
 $ECS_{strong}$ 45.0 59.0 65.2 16.0 98.2 98.8 97.0 90.2 97.8 99.0 94.6 81.4 96.0 98.6 88.0 53.6
 $ET_{weak}$ 24.0 20.4 25.2 2.0 92.2 92.4 95.2 72.4 96.2 95.6 90.0 76.2 95.0 96.4 88.0 58.0
 $ET_{medium}$ 4.6 0.6 11.6 0.0 89.8 74.0 89.6 45.0 96.0 96.0 82.2 60.8 99.0 96.0 78.0 50.2
 $ET_{strong}$ 0.4 0.0 8.4 0.0 96.4 54.8 87.4 26.8 95.2 94.8 70.0 44.8 95.0 93.0 69.0 48.0
 $LCS_{weak}$ 72.2 65.2 57.0 12.2 97.2 98.4 97.6 87.2 97.4 98.6 93.6 81.4 96.0 98.2 85.0 55.0
 $LCS_{medium}$ 44.6 34.0 17.4 1.6 97.4 97.8 97.4 68.4 97.4 99.4 88.2 75.6 97.0 94.0 75.8 51.0
 $LCS_{strong}$ 11.8 2.0 9.8 0.0 95.4 94.2 90.4 51.8 94.0 96.0 74.0 40.4 94.0 93.0 72.0 51.0
 $UB_{weak}$ 3.4 1.8 10.6 3.8 98.2 96.4 98.4 89.6 97.8 97.4 93.2 82.8 94.0 96.0 79.0 53.0
 $UB_{medium}$ 0.2 0.0 0.0 0.0 96.8 51.0 89.4 36.2 96.6 96.8 88.4 73.4 93.0 93.2 74.0 55.0
 $UB_{strong}$ 0.0 0.0 0.0 0.0 90.8 9.4 68.0 10.2 82.0 83.4 49.4 30.8 82.0 96.0 52.0 41.4
@@
Table 2: Robustness of VLA models in the simulator under various sensor attacks. LB: Laser Blinding; LP: Light Projection; ECS: EM Color Strip; ET: EM Truncation; LCS: Laser Color Strip; UB: Ultrasound Blur; VD: Voice DoS; VS: Voice Spoofing.
@@
Defense Method OpenVLA OpenVLA-OFT $\pi$0 $\pi$0-fast
@@
 AT $LB_{medium}$ 78.8 84.0 76.6 42.2 98.6 99.0 97.0 93.2 98.4 97.6 93.8 77.2 97.2 97.8 86.2 57.2
 AT $LP_{medium}$ 66.2 74.4 28.4 11.8 98.6 97.6 96.8 91.6 97.8 96.2 94.4 78.6 97.4 97.8 86.0 58.2
 AT $ECS_{medium}$ 72.0 76.0 63.6 34.8 98.2 98.0 96.8 92.4 98.6 96.4 94.4 79.6 95.6 98.6 84.2 58.0
 AT $ET_{medium}$ 35.6 3.4 23.4 0.0 99.4 96.6 97.0 79.8 98.4 97.6 92.4 76.4 96.4 97.6 82.6 52.4
 AT $LCS_{medium}$ 82.0 85.2 69.2 36.2 98.0 98.4 96.8 90.0 98.4 98.4 91.6 78.0 97.0 98.0 84.6 59.4
 AT $UB_{medium}$ 56.8 66.6 46.2 8.0 97.6 99.0 97.2 92.2 97.6 97.4 94.4 82.6 97.8 98.6 83.2 57.8
@@
Answer 1: Physical sensor attacks succeed on VLAs.
Performance without sensor attacks. We first evaluate the performance of four VLA models after fine-tuning on Libero datasets. As shown in Table 2, these models demonstrate strong performance, achieving up to 90% TSR on simple tasks such as Libero-Spatial and Libero-Object. For long-horizon tasks, OpenVLA-OFT maintains good performance. These results indicate that VLA models possess robust task execution capabilities across diverse scenarios.
Performance against simulated sensor attacks. Then, we evaluate the performance of four VLA models against simulated sensor attacks. As shown in Table 2, all VLA models exhibit vulnerability to sensor attacks. The degree of performance degradation differs depending on the specific VLA architecture, attack type, and attack intensity. In most scenarios, particularly under strong attacks or long-horizon tasks, model performance collapses catastrophically.
@@
Answer 2.1: The impacts of attacks on VLAs vary.
Camera attacks, including Laser Blinding (LB), EM Truncation (ET), and Ultrasound Blur (UB), highly disrupt the visual information of VLA models, especially under medium to strong intensity settings. The absence of object locations and class identities leads to severe task failures, resulting in unexpected actions or potentially harmful operations. In contrast, other camera attacks, such as Light Projection (LP), Laser Color Strip (LCS), and EM Color Strip (ECS), are comparatively less severe. These attacks interfere with the models’ attention by injecting perturbation rather than destroying critical visual features. Consequently, the primary objects and intended goals typically remain discernible, resulting in less degradation in TSR. Voice DoS attacks adversely affect VLA model performance, with the severity primarily determined by dataset characteristics. In the Libero-Goal dataset, where robots must execute varying instructions within fixed scenes, the absence of voice instructions leaves VLA models unable to determine appropriate actions. Conversely, datasets featuring unique scene-instruction correspondences (e.g., Libero-Spatial, Libero-Object, and Libero-Long) enable VLA models to infer correct instructions from visual context alone, thereby reducing the attack’s effectiveness. The effectiveness of the voice spoofing attack is strongly correlated with the target VLA model’s semantic understanding and instruction-following capabilities. Compared to pi0 and pi0-fast, OpenVLA and OpenVLA-OFT employ LLM backbones, making them more vulnerable to malicious instruction injection. OpenVLA-OFT incorporates a FiLM module that leverages task-specific language embeddings to modulate visual features, thereby enhancing instruction-following capabilities. Consequently, OpenVLA-OFT exhibits the most substantial performance degradation under this attack.
Answer 2.2: VLAs exhibit different robustness against sensor attacks.
@@
Attack Method OpenVLA OpenVLA-OFT $\pi$0 $\pi$0-fast
None (Baseline) 5/10 8/10 10/10 10/10
 Laser Blinding 0/10 0/10 0/10 0/10
 Light Projection 0/10 1/10 1/10 0/10
 EM Color Strip 0/10 0/10 6/10 0/10
 EM Truncation 0/10 0/10 0/10 0/10
 Laser Color Strip 0/10 0/10 3/10 0/10
 Ultrasound Blur 0/10 0/10 1/10 0/10
@@
We first evaluate the benign performance to establish baseline model availability. Subsequently, we inject malicious signals into sensors in real-world scenarios using attack parameters identified through simulation. Table 4 presents the results, demonstrating strong alignment between physical and simulation conclusions, thereby validating the effectiveness of the attack parameters searched in the simulation. As shown in Figure 5, the attacks could induce four distinct consequences: 1) Previously closed grippers unexpectedly release, causing objects to fall and sustain damage; 2) robotic arms or grippers collide with objects or environmental structures, damaging either the objects or the grippers themselves; 3) robotic arms grasp incorrect objects, preventing successful task completion; 4) robotic arms exhibit erratic movement patterns, resulting in chaos and energy waste.